Grade 12th pass9 grade science

Molecular mass of Na2​SO4​ . 10H2​O is:

Harshit Singh
4 Years agoGrade 12th pass

1 Answer

Yash Patari
4 Years ago
Dear student

Molecular mass of Washing Soda or Na2SO4•10H20= 23×2+32+16×4+10(1×2+16)

=46+32+64+180

=78+244

=322

thanks and reggards

teamm askiitians